Journalist Report – March 06th

Crew 223 Journalist Report 06Mar2020
Author: Clément Plagne, Crew Journalist
Sol 5
Title: Grounded
Mars is a colourful place. Today on EVA, we explored the North Ridge, a
hilly area a few miles from the Station. The exploration was rather
athletic, and we once again saw all sorts of terrain as we trudged up
and down the steep sandy slopes. It felt like we walked for hours and
miles, despite the area being rather small and us being done in about an
hour. It probably was because the terrain changes so much – in a matter
of minutes we went from red hills to rocky areas to grey, Moon-like
zones. It’s also probably because the heavy suits make us slower. Most
certainly, it’s a bit of both. The other great news is that the more
comfortable suits that had failed on our first EVAs have been handsomely
fixed by Luc and Aurélien, and now work fully. Begone, back pains!
The ground out there is barren and lifeless. Still, there’s somewhere
not too far where the soil is life-giving and fruitful. The greenhouse
is in full bloom thanks to the good work of Valentin, our devoted
GreenHab officer. Aromatics, radishes, carrots and many others are busy
growing all day long and may at some point be food for us. But there’s
another, odder, thing growing in the warmth of the GreenHab. In glass
tanks lives a green, bubbly mass called spirulina. They’re algae that
shine by the low area needed to grow it, and the massive amounts of
nutrition value it creates with very little input. In those two tanks
that occupy very little of the water consumption and space of the
greenhouse, they have the potential of feeding us much better than all
the rest. It’s not as tasty as rosemary or basil, but some day we may be
forced to be as efficient as possible, and we’ll be happy to have it.
The Science Dome is also on the cutting edge of vegetal research. Today
was the official start of the Music for Plants experiment. It posits
that in harsh environments, sound waves may influence the durability of
plants – in other words, plants may enjoy music! To test this, Valentin
puts different batches of radishes and watercress under strong UV
lights, basically giving them sunburns, and tests different sounds on
the plants in the meantime, seeing how they evolve every day. With the
thin atmosphere of Mars, plants will inevitably be put under stress from
sunrays, so finding out how best to protect them and have an agriculture
on another planet is fundamental. In that same Science Dome, a tower
breathes. No soil this time: little pods on the tower are filled with a
special foam and fed with a nutritive liquid mix, and plants grow just
as well. Aeroponics and vertical culture are possible keys to
efficiently feed a colony on Mars, and working with the first
small-scale examples of these is a privilege for us.

Astronomy Report – March 06th

  
Crew 223 Astronomy Report Mar 06 2020

Name: Florian Delpech
Crew: 223
Date: 3/6/2020

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: The last 6 galaxies for the research
project: NGC 4579 ; NGC 4636 ; NGC 4621 ; NGC 4666 ; NGC 5047 ; NGC
3992. I am happy to see that I am lucky with the weather!

Images submitted with this report: 4 very different galaxies where
supernovae could maybe be fine (they have to be compared to the
reference image and the ones taken during Eishi's mission).
I am impressed to see how diverse the galaxies are!

Problems Encountered: The calibration file MDRS14_mfalt_V is missing, I
used an older one

MUSK OBSERVATORY

Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Sol Summary – March 06th

Crew 223 Sol Summary Report 06-03-2020

Sol: 5 Summary Title: Mid rotation is coming to fast

Author’s name: Aurélien Mure – Crew Commander

Mission Status: Everything’s fine

Sol Activity Summary:

The crew woke up at 6:50am. The sport session began at 7:00am. Our
training is made of different exercises repeated 3 times 20s on Monday
and Tuesday, twice 30s on Wednesday and Thursday and once 1min on Friday
and Saturday.
The one minute duration exercise is really exhausting for the muscles
and it was a little bit painful this morning!
During the EVA, the team has changed the batteries of the LOAC and
MegaAres and then explored North Ridge. They were truly happy about what
they saw considering their smile at the end. I think our communications
are getting better and better in EVA. That is why we worked with Luc on
a protocole enabling us to perform a non verbal communication EVA.
Florian has worked on his astronomy researches for a while. Then I
worked with Florian and Blandine on the solar oven that is ready to go
our tomorrow.
Clément and Valentin took care of some cress and radishes for our Music
for plants experiment.
Marion tried to unlock the code for our display panel.
Tonight, we realise that midrotation is arriving too fast. For me it is
a real difference as I was involved in a 3 weeks rotation last year, it
changes a lot!

Look Ahead Plan:

The non verbal EVA planned is really exciting, and the solar oven will
be set up

Anomalies in work: None.

Weather: Cloudy and Sunny alternatively, -2 at 7:00

Crew Physical Status: All crew members safe and sound.

EVA: LOAC and MegaAres batteries changed North Ridge exploration

Reports to be filed: Journalist Report, Operations report, EVA Request,
Green Hab report, EVA Report, Astronomy report

Support Requested: None

GreenHab Report – March 06th

  

Crew 223 GreenHab Report - 06-03-2020 GreenHab Officer: Valentin
BELLEMIN-LAPONNAZ

Environmental control: Ambient with Door open

Average temperature: 22.5°C
Low temperature: 15°C
High temperature: 25°C

Hours of supplemental light: 4 hours (from 9P.M to 1 a.m.)

Daily water usage for crops: 12.6 gallons

Daily water usage for research and/or other purposes: 0 gallon

Water in Blue Tank— 153.7 gallons

Time(s) of watering for crops: 3 (1 only for tomatoes and radish)

Changes to crops: NA

Narrative: This morning, Aurelien MURE watered all the plants, as I were
on EVA. After lunch, I carried out the experience Music For Plants. I
decided to use both young and advanced radish plants. I mounted the UV
chamber and the speaker and did the first UV incubation and listening
for the batch of plants. For dinner, I harvested basil, oregano and
cilantro.

Harvest: 3g Basil / 3g Cilantro / 4g oregano

Support/supplies needed: NA 

EVA Report – March 06th

Crew 223 EVA Report 06-03-2020
EVA #5
Author: Clément Plagne
Purpose of EVA: Change the batteries for MegaARES and LOAC.
Familiarising ourselves with the terrain and the hiking required for
exploration of the environment. Testing the recently repaired suits.
Start Time: 09:07
End Time: 11:04
Narrative:
We left the airlock at 09:12, a bit later than planned. Checks for rover
status and water level went nominally. Valentin was tasked with testing
the Honda ATV, that had failed to start previously – attempt was a
success and Valentin reached first gear to confirm ATV was running.
We reached the LOAC/MegaARES area at 09:20, and battery replacement
tasked went nominally as usual.
We took off for North Ridge and parked the ATVs near Stream Bed
Connector, heading west by foot. We witnessed a helicopter making a
close approach near us and kept going. We hiked through steep terrain,
trying to make our way to the top of the Ridge but stopping due to
safety concerns.
Our hike lasted for a bit less than an hour, going from hilly, red-hued
hills to rockier zones, to gray, nearly Moon-like zones. We decided to
leave earlier instead of attempting a tough looking climb in a rocky area.
We arrived back near the hab at 10:56, and were back in the airlock by
11:04. All suits worked perfectly.
Destination: North Ridge.
Coordinates (use UTM NAD27 CONUS): 12S 518500 4250800 and 12S 518500 425000
EVA Participants: Clément Plagne (EVA Leader, Journalist), Blandine
Gorce (HSO), Valentin Bellemin-Laponnaz (GreenHab Officer), Florian
Delpech (Astronomer)
Road(s) and routes per MDRS Map: Cow Dung Road
Mode of travel: Driving (and walking when off-road)
Vehicles you will be using (If applicable): Rovers Deimos/Perseverance
and Curiosity

Operations Report – March 06th

  

Crew 223 Operations Report 06/03/2020

SOL: 5

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Thursday generator:
Off at 4:30am
On at 9:10pm SOC was 84%

Friday generator:
Off at 6:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 75%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149721.59 units

Water (static tank) – 307 gallons

Static to Loft Pump used – yes

Water in GreenHab – 153.7 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: yes
Hours: 149.7
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: no
Hours: 133.7
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.5
Beginning charge: 100%
Ending charge: 80%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Honda started during the EVA this morning, and we had no
problem with it. I started the 4 ATVs during the engineer EVA for
battery check.
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: Nothing to report

HabCar used and why, where? The HabCar is parked at the MDRS Campus.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA.
No problem.

Summary of Hab operations: One of the gaz button of the gaz stove is
broken (the one for the upper left).

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Florian built the solar hoven.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report